Exchange Corp. is a company that acts as a facilitator in tax-favored real estate swaps. Such swaps, know as 1031 exchanges, permit participants to avoid some or all of the capital gains taxes that would otherwise be due. The bookkeeper for the company has been asked to prepare a report for the company to help its owner/manager analyze performance. The first such report appears below:

  

Exchange Corp
Analysis of Revenues and Costs
For the Month Ended May 31
Actual
Unit Revenues
and Costs
  Planning Budget
Unit Revenues
and Costs
Variances
  Exchanges completed 30     25   
  Revenue $ 580     $ 670    $ 90   U
  Expenses:
    Legal and search fees 233    215    18   U
    Office expenses 155     245    90   F
    Equipment depreciation 25     30    5   F
    Rent 75     90    15   F
    Insurance 15     18    3   F
  Total expense 503     598    95   F
  Net operating income $ 77     $ 72    $ 5   F

  

Note that the revenues and costs in the above report are unit revenues and costs. For example, the average office expense is $245 per exchange completed on the planning budget; whereas, the average actual office expense is $155 per exchange completed.

     Legal and search fees is a variable cost; office expenses is a mixed cost; and equipment depreciation, rent, and insurance are fixed costs. In the planning budget, the fixed component of office expenses was $4,850.

     All of the company’s revenues come from fees collected when an exchange is completed.

  

Required:
1. Is the report prepared by the bookkeeper useful as a performance report?
  • Yes

  • No

  

2.

Complete a performance report that would help the owner/manager assess the performance of the company in May. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)

Exchange Corp
Flexible Budget Performance Report
For the Month Ended May 31
Actual Results Revenue and Spending Variances Flexible Budget Activity Variances Planning Budget
Exchanges completed
Revenue
Expenses:
Legal and search fees
Office expenses
Equipment depreciation
Rent
Insurance
Total expense
Net operating income
Answer #1
1
NO, the report prepared by bookkeeper is not a useful performance report
2
Actual results Revenue and spending variances Flexible budget Activity variances Planning budget
Exchanges completed 30 30 25
Revenue 17400 2700 U 20100 3350 F 16750
Expenses:
Legal and search fees 6990 540 U 6450 1075 U 5375
Office expenses 4650 1730 F 6380 255 U 6125
Equipment Depreciation 750 0 None 750 0 None 750
Rent 2250 0 None 2250 0 None 2250
Insurance 450 0 None 450 0 None 450
Total expense 15090 1190 F 16280 1330 U 14950
Net operating income 2310 1510 U 3820 2020 F 1800
Office expenses:
Variable component = (6125-4850)/25 = $51
Office expenses in flexible budget = 4850+(30*51) = $6380